OpenStudy (anonymous):

Please Help!!! (3/2b^4)^3

OpenStudy (mathstudent55):

Is this the question? \(\LARGE \left( \dfrac{3}{2b^4} \right)^3\)

OpenStudy (mathstudent55):

When you raise a fraction to a power, you raise the numerator and the denominator to that power.

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

First, raise the numerator and the denominator to the 3rd power: \(\LARGE \left( \dfrac{3}{2b^4} \right)^3 = \dfrac{3^3}{(2b^4)^3}\)

OpenStudy (mathstudent55):

The numerator is simple, \(3^3\) is simply \(27\).

OpenStudy (mathstudent55):

Now we need to work on the denominator.

OpenStudy (mathstudent55):

When you raise a product to a power, you raise each factor of the product to that power.

OpenStudy (mathstudent55):

\( \LARGE =\dfrac{27}{2^3(b^4)^3}\)

OpenStudy (mathstudent55):

\(2^3\) is smple. It is \(8\).

OpenStudy (mathstudent55):

Now we need to deal with \((b^4)^3\).

OpenStudy (mathstudent55):

When you raise a power to a power, you multiply powers.

OpenStudy (mathstudent55):

\( \LARGE =\dfrac{27}{8b^{12}}\)

OpenStudy (anonymous):

that would be b12?

OpenStudy (anonymous):

Thank you soooo much!!!!!!! I needed that help thank you

OpenStudy (mathstudent55):

b to the 12th power.

OpenStudy (mathstudent55):

Here are the three rules of exponents we used: \(\left(\dfrac{a}{b}\right)^n = \dfrac{a^n}{b^n}\) \((ab)^n = a^nb^n\) \((a^m)^n = a ^{mn} \)

OpenStudy (mathstudent55):

You are welcome.